The physical plant at the main campus of a large state university recieves daily requests to replace florecent lightbulbs. The distribution of the number of daily requests is bell-shaped and has a mean of 57 and a standard deviation of 7. Using the 68-95-99.7 rule, what is the approximate percentage of lightbulb replacement requests numbering between 57 and 78?
Answer #1

= 57

= 7

To find P(57 < X < 78):

Z = (78 - 57)/7 = 3

By 68 - 95- 99.7 Rule, Percentage of data points within 3 Standard Deviation from mean = 99.7 %

Since here only mid value to3 standard deviation on RHS is required,:

Percentage of light bulb replacement requests numbering between 57 and 78 = 99.7/2 = 49.85 %

So,

Answer is:

49.85%